OpenStudy (anonymous):

Look at the formula for the volume of a cylinder. V = r2h Which is a solution for h?

OpenStudy (anonymous):

Look at the formula for the volume of a cylinder. V = pi r2h Which is a solution for h?

OpenStudy (anonymous):

h=v/2r

OpenStudy (anonymous):

\[ h = \frac{V}{2\pi r} \]

OpenStudy (anonymous):

check your formula, yopu missed pi